### v1

@@ Abstract
We introduce Paper2Agent, an automated framework that converts research papers into AI agents. Paper2Agent transforms research output from passive artifacts into active systems that can accelerate downstream use, adoption, and discovery. Conventional research papers require readers to invest substantial effort to understand and adapt a paper’s code, data, and methods to their own work, creating barriers to dissemination and reuse. Paper2Agent addresses this challenge by automatically converting a paper into an AI agent that acts as a knowledgeable research assistant. It systematically analyzes the paper and the associated codebase using multiple agents to construct a Model Context Protocol (MCP) server, then iteratively generates and runs tests to refine and robustify the resulting MCP. These paper MCPs can then be flexibly connected to a chat agent (e.g. Claude Code) to carry out complex scientific queries through natural language while invoking tools and workflows from the original paper. We demonstrate Paper2Agent’s effectiveness in creating reliable and capable paper agents through in-depth case studies. Paper2Agent created an agent that leverages AlphaGenome to interpret genomic variants and agents based on ScanPy and TISSUE to carry out single-cell and spatial transcriptomics analyses. We validate that these paper agents can reproduce the original paper’s results and can correctly carry out novel user queries. By turning static papers into dynamic, interactive AI agents, Paper2Agent introduces a new paradigm for knowledge dissemination and a foundation for the collaborative ecosystem of AI co-scientists.
@@ 1 Introduction
Recent advances highlight the promise of agents for accelerating discovery. For example, the Virtual Lab framework organizes teams of AI scientist agents that collaboratively design and execute research projects across biology and chemistry [6]. Similarly, Google’s AI co-scientist serves as a virtual collaborator, assisting with hypothesis generation and research proposal development [7]. Sakana AI’s co-scientist aims for automation of the research lifecycle—from ideation to publication [8]. FutureHouse provides an AI scientist platform designed for diverse scientific tasks [9]. Alongside these general-purpose platforms, specialized agents are also emerging for specific domains [10]. For example, CellVoyager introduces an agentic system for autonomous analysis of single-cell omics data [11]. Biomni is an AI agent for diverse biological tasks [12]. These systems demonstrate that agents can not only execute code, but also generate hypotheses, evaluate uncertainty, and adapt methods to new datasets. Paper2Agent complements this emerging paradigm by generalizing the concept: any research paper can be converted into an agent that embodies the knowledge and methods described in the publication.
@@ 2 Results
Figure 2: Overview of the Paper2Agent-generated AlphaGenome agent. (A) Construction of the AlphaGenome MCP server and agent. (B) Exposed MCP tools and resources enabling variant scoring and visualization. (C) Benchmarking shows 100% accuracy on queries for reproduction and new tasks. (D) Automated planning and interpretation of GWAS loci through iterative planning–action–observation cycles by the AlphaGenome agent.
@@
Importantly, the tools generated by Paper2Agent are designed with flexible, well-annotated input parameters. For example, the visualize_variant_effects() tool exposes a rich set of options that make it adaptable to diverse use cases (Figure  2B). Given an input genetic variant, the AlphaGenome agent can select the organism to analyze (human or mouse), adjust the sequence context length around the variant, toggle different modalities—such as RNA-seq, ATAC-seq, or ChIP-seq histone tracks. Moreover, each MCP tool embeds a traceable link to the original GitHub source code, ensuring transparency and reproducibility. By connecting an AI agent with the AlphaGenome MCP, the system creates the AlphaGenome agent.
@@
Next, we benchmarked the Paper2Agent-generated AlphaGenome agent in producing numerical results and figures relative to human experts configuring and running the code manually (Figure  2C). We manually curated 15 example queries directly from the AlphaGenome tutorial, such as "Score variant chr3:58394738:A>T using ATAC-seq predictions for motor neuron cells (CL:0000100). What is the quantile_score for this cell type?", "Make DNase-seq predictions for sequence ’GATTACA’ (padded to 2048 length) for lung tissue (UBERON:0002048). What is the nonzero_mean value in the dnase metadata". The AlphaGenome agent achieved 100% accuracy on these queries, precisely matching all reported values. To assess generalizability and guard against potential overfitting to the original examples, we also manually curated a set of novel queries that were not present in either the paper or its codebase (see Supplementary Table 2). These included previously untested variant positions, allelic substitutions, and tissue–cell type contexts, such as "Analyze variant chr9:98765432:T>C with DNASE predictions for muscle cells (CL:0000187). What is the quantile_score for muscle tissue?" and "Analyze histone ChIP-seq metadata for neuronal stem cells. What is the nonzero_mean value for H3K4me3 in neuronal stem cells (CL:0000100)?" The AlphaGenome agent again achieved 100% accuracy, faithfully producing the expected numerical outputs, which we verified by manual execution of the original AlphaGenome code.
@@
Finally, we demonstrated that the AlphaGenome agent enables automatic interpretation of GWAS loci and validation of the analysis in the original paper. We considered the example of interpreting why the genetic variant chr1:109274968:G>T is associated with low-density lipoprotein cholesterol that was presented in the original AlphaGenome paper (Figure  2D). Based on the tools available, the AlphaGenome agent constructs a step-by-step plan to solve this task. This plan includes generating input files, scoring variants across multiple modalities, filtering results for trait-relevant tissues, creating modality-specific visualizations (chromatin accessibility, histone marks, transcription factor binding, and splicing), and assembling a comprehensive interpretation report. The agent then executes these actions using implemented tools, such as score_variant() and visualize_tf_binding(), automatically refining its strategy through iterative observation and feedback. A final report is then presented to provide a unified interpretation of the regulatory impact of the variant, integrating evidence across modalities and tissues.
@@
Interestingly, the AlphaGenome agent prioritizes SORT1 as the most likely causal gene, whereas the original paper emphasized CELSR2 and PSRC1. The agent favors SORT1 for two reasons: 1) a high quantile score (0.99982) indicating a strong predicted impact on SORT1 expression in liver tissue. Here, the quantile score reflects how extreme the variant’s predicted effect relative to other variants 2) SORT1 encodes sortilin, directly involved in LDL/VLDL secretion [20]. We manually queried the GTEx [21] eQTL data and confirmed that this variant is a significant liver eQTL for SORT1 (p = 1.1e-65) in liver. However, both CELSR2 and PSRC1 also exhibit high AlphaGenome quantile scores (0.99998 each) and significant eQTL associations in GTEx liver (p = 4.7e-46 and 8.5e-50, respectively). This shows the inherent difficulty in confidently assigning causal genes at complex GWAS loci where the variants are eQTLs for multiple nearby genes [22, 23].
@@
We next present the Paper2Agent-generated paper agent for TISSUE [18], a recent paper that developed a new method for uncertainty-aware single-cell spatial transcriptomics analysis (Figure 3A). This case study reflects a common scenario: a new methodology paper is published, and researchers want to apply the method to their own data but lack the time to navigate the codebase, configure the environment, and grasp the method’s features and input requirements. Paper2Agent addresses these challenges by automatically generating ready-to-use agents for diverse papers and providing Q&A support to guide input preparation and clarify what the method can do.
@@ 3 Discussion
Our initial focus has been on methodological papers, since they offer the clearest use case. Such papers typically describe algorithms, models, or computational workflows that other researchers seek to adopt, but whose deployment often requires substantial technical expertise. Converting them into agents allows the methods to be applied to new problems without the overhead of mastering complex software ecosystems. In future work, we plan to extend Paper2Agent to other forms of research output, including data resources and discovery papers. In those contexts, the agent’s role may shift from computation to interpretation, curation, or explanation, guiding users through datasets or contextualizing new insights for diverse scientific communities.
@@
To better quantify this ease of reproducibility and agentification, we have introduced a benchmarking approach based on manually evaluated examples from the paper as well as novel examples meant to test generalizability. With this approach, we showed, for example, that the AlphaGenome agent was able to execute both tutorial-based and novel queries with 100% accuracy. This approach, however, is limited by expert knowledge of the paper and method and manual implementation and review. A future direction is to further streamline this process with additional agentic frameworks, e.g. with LLM-as-judge evaluations [24].
Another consideration is the scale of agentification. While the paper is the conventional unit of scientific communication, it is not always the best unit for agentification. In many fields, an idea evolves across a sequence of publications, each adding refinements, benchmarks, or applications. In such cases, the most useful agent may not represent a single paper but rather a collection of related works aggregated into a coherent interface. A single MCP can encapsulate multiple related papers. We plan to work on extensions of Paper2Agent to flexibly accommodate this broader scope.
@@
Finally, once scientific knowledge is encoded in active agents rather than static artifacts, the potential extends beyond individual use. Agents could interact with one another, linking methods to datasets or combining insights from different domains. For example, an agent representing a new analytical method could collaborate with an agent representing a newly released dataset, jointly producing analyses that neither artifact could support alone. Communities of such agents could form a dynamic, interoperable layer of scientific intelligence, accelerating connections across disciplines. Paper2Agent thus points toward a future in which scientific communication is not only about describing results, but also about creating interactive, collaborative entities that embody and extend the research.

### v2

@@ Abstract
We introduce Paper2Agent, an automated framework that converts research papers into AI agents. Paper2Agent transforms research output from passive artifacts into active systems that can accelerate downstream use, adoption, and discovery. Conventional research papers require readers to invest substantial effort to understand and adapt a paper’s code, data, and methods to their own work, creating barriers to dissemination and reuse. Paper2Agent addresses this challenge by automatically converting a paper into an AI agent that acts as a knowledgeable research assistant. It systematically analyzes the paper and the associated codebase using multiple agents to construct a Model Context Protocol (MCP) server, then iteratively generates and runs tests to refine and robustify the resulting MCP. These paper MCPs can then be flexibly connected to a chat agent (e.g. Claude Code) to carry out complex scientific queries through natural language while invoking tools and workflows from the original paper. We demonstrate Paper2Agent’s effectiveness in creating reliable and capable paper agents through in-depth case studies. Paper2Agent created an agent that leverages AlphaGenome to interpret genomic variants and agents based on ScanPy and TISSUE to carry out single-cell and spatial transcriptomics analyses. We validate that these paper agents can reproduce the original paper’s results and can correctly carry out novel user queries. Paper2Agent automatically created AI co-scientist that identified new splicing variant associated with ADHD risk. By turning static papers into dynamic, interactive AI agents, Paper2Agent introduces a new paradigm for knowledge dissemination and a foundation for the collaborative ecosystem of AI co-scientists.
@@ 1 Introduction
Recent advances highlight the promise of agents for accelerating discovery. For example, the Virtual Lab framework organizes teams of AI scientist agents that collaboratively design and execute research projects across biology and chemistry [14]. Similarly, Google’s AI co-scientist serves as a virtual collaborator, assisting with hypothesis generation and research proposal development [16]. Sakana AI’s co-scientist aims for automation of the research lifecycle—from ideation to publication [23]. FutureHouse provides an AI scientist platform designed for diverse scientific tasks [17]. Alongside these general-purpose platforms, specialized agents are also emerging for specific domains [18]. For example, CellVoyager introduces an agentic system for autonomous analysis of single-cell omics data [19]. Biomni is an AI agent for diverse biological tasks [20]. These systems demonstrate that agents can not only execute code, but also generate hypotheses, evaluate uncertainty, and adapt methods to new datasets. Recent work has also explored automatic code generation from scientific text [27, 26]. Paper2Agent complements this emerging paradigm by generalizing the concept: any research paper can be converted into an agent that embodies the knowledge and methods described in the publication.
@@ 2 Results
Figure 2: Overview of the Paper2Agent-generated AlphaGenome agent. (A) Construction of the AlphaGenome MCP server and agent. (B) Benchmarking the AlphaGenome agent on tutorial-based and novel queries shows 100% accuracy, above that achieved by the general agents Claude with the AlphaGenome codebase (Claude + Repo) or Biomni. (C) The AlphaGenome agent also shows improved computational efficiency in terms of reduced run-time for each query compared to Claude + Repo and Biomni agents. (D) Automated planning and interpretation of GWAS loci through iterative planning–action–observation cycles by the AlphaGenome agent.
@@
Importantly, the tools generated by Paper2Agent are designed with flexible, well-annotated input parameters. For example, the visualize_variant_effects() tool exposes a rich set of options that make it adaptable to diverse use cases (Supplementary Figure 1). Given an input genetic variant, the AlphaGenome agent can select the organism to analyze (human or mouse), adjust the sequence context length around the variant, toggle different modalities—such as RNA-seq, ATAC-seq, or ChIP-seq histone tracks. Moreover, each MCP tool embeds a traceable link to the original GitHub source code, ensuring transparency and reproducibility. By connecting an AI agent with the AlphaGenome MCP, the system creates the AlphaGenome agent.
@@
Next, we benchmarked the Paper2Agent-generated AlphaGenome agent in producing numerical and qualitative results and figures relative to human experts configuring and running the code manually (Figure 2B). We also compared the AlphaGenome agent to two other agentic systems: 1) Claude Code with access to the AlphaGenome repo (referred to as Claude + Repo) and 2) Biomni. We manually curated 15 example queries directly from the AlphaGenome tutorial, such as "Score variant chr3:58394738:A>T using ATAC-seq predictions for motor neuron cells (CL:0000100). What is the quantile_score for this cell type?", "Make DNase-seq predictions for sequence ’GATTACA’ (padded to 2048 length) for lung tissue (UBERON:0002048). What is the nonzero_mean value in the dnase metadata". The AlphaGenome agent achieved 100.0% (15/15) accuracy on these queries, higher than Claude + Repo at 60.0% (9/15), and Biomni at 40.0% (6/15).
To assess generalizability and guard against potential overfitting to the original examples, we also manually curated a set of novel queries that were not present in either the paper or its codebase. These included previously untested variant positions, allelic substitutions, and tissue–cell type contexts, such as "Analyze variant chr9:98765432:T>C with DNASE predictions for muscle cells (CL:0000187). What is the quantile_score for muscle tissue?" and "Analyze histone ChIP-seq metadata for neuronal stem cells. What is the nonzero_mean value for H3K4me3 in neuronal stem cells (CL:0000100)?" The AlphaGenome agent again achieved 100.0% (15/15) accuracy, as verified by manual execution of the original AlphaGenome code. Claude + Repo achieved the next highest accuracy at 80.0% (12/15) and Biomni solved 60.0% (9/15) of the queries correctly.
The AlphaGenome agent also consistently outperformed both agent systems in terms of compute efficiency, showing a median decrease in run time for the tutorial-based benchmark of 1.8x and 3.1x compared to Claude + Repo and Biomni, respectively (Figure 2C). For the novel benchmark, we observed a median run time improvement of 3.2x and 4.6x, respectively. These results highlight the improved reliability and efficiency achieved by Paper2Agent-generated agents compared to other agentic systems.
@@
Finally, we demonstrated that the AlphaGenome agent enables automatic interpretation of Genome-Wide Association Study (GWAS) loci and validation of the analysis in the original paper. We considered the example of interpreting why the genetic variant chr1:109274968:G>T is associated with low-density lipoprotein cholesterol that was presented in the original AlphaGenome paper (Figure 2D). Based on the tools available, the AlphaGenome agent constructs a step-by-step plan to solve this task. This plan includes generating input files, scoring variants across multiple modalities, filtering results for trait-relevant tissues, creating modality-specific visualizations (chromatin accessibility, histone marks, transcription factor binding, and splicing), and assembling a comprehensive interpretation report. The agent then executes these actions using implemented tools, such as score_variant() and visualize_tf_binding(), automatically refining its strategy through iterative observation and feedback. A final report is then presented to provide a unified interpretation of the regulatory impact of the variant, integrating evidence across modalities and tissues.
@@
Interestingly, the AlphaGenome agent prioritizes SORT1 as the most likely causal gene, whereas the original paper emphasized CELSR2 and PSRC1. The agent favors SORT1 for two reasons: 1) a high quantile score (0.99982) indicating a strong predicted impact on SORT1 expression in liver tissue. Here, the quantile score reflects how extreme the variant’s predicted effect relative to other variants 2) SORT1 encodes sortilin, directly involved in LDL/VLDL secretion [5]. We manually queried the GTEx [6] eQTL data and confirmed that this variant is a significant eQTL for SORT1 (p = 1.1e-65) in liver. However, both CELSR2 and PSRC1 also exhibit high AlphaGenome quantile scores (0.99998 each) and significant eQTL associations in GTEx liver (p = 4.7e-46 and 8.5e-50, respectively). This result shows the inherent difficulty in confidently assigning causal genes at complex GWAS loci where the variants are eQTLs for multiple nearby genes [7, 11].
@@
We next present the Paper2Agent-generated agent for TISSUE [10], a recent paper that developed a new method for uncertainty-aware single-cell spatial transcriptomics analysis (Figure 3A). This case study reflects a common scenario: a new methodology paper is published, and researchers want to apply the method to their own data but lack the time to navigate the codebase, configure the environment, and grasp the method’s features and input requirements. Paper2Agent addresses these challenges by automatically generating ready-to-use agents for diverse papers and providing Q&A support to guide input preparation and clarify what the method can do.
@@
Paper2Agent enables autonomous AI-driven collaboration and discovery
Human scientific collaboration often advances by combining insights from multiple, disparate papers—for instance, when a newly developed method is applied to a recently published dataset to generate fresh discoveries. However, this process is typically slow and labor-intensive. Paper2Agent enables a new mode of AI-driven collaboration in which AI paper agents can interact directly with each other. The agent of a new method paper can autonomously collaborate with the agent of a new data paper to perform analyses, test hypotheses, and generate new insights.
As an illustrative case study, we consider the AlphaGenome method paper [13] and a recent GWAS of Attention-Deficit/Hyperactivity Disorder (ADHD) data and discovery paper [28], representing a common scenario where a new method and a new dataset/discovery become available. Researchers may want to collaborate to integrate AlphaGenome with the data and insights from the GWAS study, but doing this manually can take weeks.
Paper2Agent automatically creates MCPs for AlphaGenome and the ADHD GWAS data. The data agent in Paper2Agent automatically converts the paper-associated datasets—such as released GWAS summary statistics and supplementary tables reporting genetic discoveries—into standardized MCP resources, enabling seamless integration with analytical agents for downstream analyses.
These MCPs are then automatically connected with a downstream AI co-scientist. Here we use Claude Code as our agentic system. This Paper2Agent-generated AI co-scientist autonomously generates new hypotheses, and designs and implements analyses to explore each hypothesis. The AI co-scientist proposed several interesting hypotheses to explore, including: 1) ADHD risk variants alter regulatory activity in brain-specific cell types. 2) AlphaGenome prioritizes causal variants within ADHD fine-mapping credible sets. 3) ADHD-associated variants disrupt transcription factor binding at FOXP family gene loci. Users can interact with the AI co-scientist in real-time, guiding it to execute the research plan for hypothesis (2) and to explore the mechanisms of the identified causal variant. Combining insights from the AlphaGenome and ADHD GWAS papers, the AI co-scientist prioritizes a single causal variant from 209 candidates identified in the fine-mapping credible sets and elucidates its molecular mechanisms contributing to ADHD risk. The agent showed that the intronic variant rs1626703 is predicted to alter MPHOSPH9 splicing and expression specifically in glutamatergic neurons (AlphaGenome quantile scores: splice junction = 1.000; RNA-seq = 0.963). AlphaGenome indicates that rs1626703 promotes exon inclusion, as evidenced by increased read coverage and strengthened splice junctions in the alternate allele, resulting in elevated MPHOSPH9 expression. MPHOSPH9 encodes an M-phase–associated phosphoprotein that localizes to centrosomes/centrioles and has been implicated in the recruitment of the CP110-CEP97 complex during cell division and ciliogenesis. [3].
Furthermore, the AI co-scientist autonomously uses AlphaGenome to prioritize causal variants across all 39 loci, completing the analysis within two hours. It automatically constructs a systematic workflow that extracts credible-set variants, performs AlphaGenome functional scoring for those variants in glutamatergic neurons, filters for protein-coding genes, and ranks variants by their maximum quantile impact scores across different functional modalities. For each locus, the AI co-scientist identifies the top variant, maps it to its target gene, summarizes its molecular effects, and compiles a comprehensive markdown report detailing the functional consequences and biological significance of the prioritized gene. The AI-generated reports for the causal variants and their mechanisms aross all 39 loci are provided in Supplementary Table 1. This workflow enables scalable hypothesis generation and mechanistic interpretation of GWAS loci in hours rather than weeks of manual review. While these hypotheses benefit from subsequent human evaluation, the workflow shifts scientific effort from manual execution to the synthesis of actionable biological insights.
Together, these results exemplify a new collaborative paradigm in which human scientists use Paper2Agent to design their own AI co-scientists and jointly formulate high-level scientific hypotheses, while the AI co-scientists autonomously execute and interpret complex analytical tasks.
Figure 5: Paper2Agent enables autonomous AI-driven collaboration and genomic discovery. (A) Paper2Agent transforms scientific papers into Model Context Protocol (MCP) resources for both methods and data, allowing AI co-scientist to integrate them and autonomously generate novel hypotheses and actionable research plans. (B) Using the ADHD GWAS dataset and the AlphaGenome method MCPs, the agent autonomously generates and tests scientific hypotheses, identifies causal variants, and interprets molecular mechanisms. The agent identified rs1626703 as a likely causal variant among 209 candidate variants. The agent then showed computationally using AlphaGenome that rs1626703 alters splicing of MPHOSPH9 and increases its expression in glutamatergic neurons, revealing a plausible causal mechanism for ADHD risk.
@@ 3 Discussion
Our initial focus has been on methodological papers, since they offer the clearest use case. Such papers typically describe algorithms, models, or computational workflows that other researchers seek to adopt, but whose deployment often requires substantial technical expertise. Converting them into agents allows the methods to be applied to new problems without the overhead of mastering complex software ecosystems. In future work, we plan to further expand Paper2Agent to other forms of research output, including data resources and discovery papers. In those contexts, the agent’s role may shift from computation to interpretation, curation, or explanation, guiding users through datasets or contextualizing new insights for diverse scientific communities.
@@
To better quantify this ease of reproducibility and agentification, we used a benchmarking approach based on human expert-evaluated examples from the paper as well as novel examples meant to test generalizability. With this approach, we showed, for example, that the AlphaGenome agent was able to execute both tutorial-based and novel queries with 100.0% accuracy, much higher than the accuracy achieved by other agents like Claude Code and Biomni. Moreover, the agent created by Paper2Agent is substantially more efficient, achieving median run time speedups of over 1.7x compared to Biomni or using Claude Code direct on the repo. Our benchmarking approach, however, is limited by expert knowledge of the paper and method and manual implementation and review. A future direction is to further streamline this process with additional agentic frameworks, e.g. with LLM-as-judge evaluations [22].
Another consideration is the scope of agentification. While the paper is the conventional unit of scientific communication, it is not always the best unit for agentification. In many fields, an idea evolves across a sequence of publications, each adding refinements, benchmarks, or applications. In such cases, the most useful agent may not represent a single paper but rather a collection of related works aggregated into a coherent interface. A single MCP can encapsulate multiple related papers. We plan to extend Paper2Agent to flexibly accommodate this broader scope.
@@
Finally, once scientific knowledge is encoded in active agents rather than static artifacts, the potential extends beyond individual use. Agents could interact with one another, linking methods to datasets or combining insights from different domains, as illustrated in our AlphaGenome-ADHD case study. Communities of such agents could form a dynamic layer of scientific intelligence, accelerating connections across disciplines and enabling a new form of AI-driven collaboration. Paper2Agent thus points toward a future in which scientific communication is not only about describing results, but also about creating interactive, collaborative entities that embody and extend the research.
@@ Acknowledgments
Supplementary Figure
Figure 6: Supplementary Figure 1. Exposed MCP tools and resources enabling variant scoring and visualization.